# BayeSN and SALT: A Comparison of Dust Inference Across SN Ia Light-curve Models with DES5YR

### v2

@@ 7 Discussion and Conclusions
Previous analyses of the host galaxy dust properties of SNe Ia using SALT and BayeSN have generally found conflicting results. SALT analyses have found large differences in the mean of the $R_{V}$ distribution, $\mu_{R_{V}}$, between SNe Ia on either side of the mass step, finding $\Delta\mu_{R_{V}}\sim 1.2-1.6$; although this approach did not, until recently, allow for achromatic intrinsic differences. BayeSN, on the other hand, has supported the existence of an intrinsic contribution to the mass step and predicted $\Delta\mu_{R_{V}}$ values which are smaller (G24) or consistent with zero (Thorp et al., 2021; Thorp & Mandel, 2022).
In this work we have for the first time carried out cross validation of the SALT and BayeSN models by applying BayeSN for inference of dust properties on SALT-based simulations similar to those used in the DES5YR cosmology analysis. Despite the fundamental differences between the models, we find that BayeSN is able to accurately recover simulated dust properties from these SALT-based simulations for a wide variety of permutations of different simulations.
Having established this consistency based on simulations, we then apply BayeSN to real data from the DES5YR sample. We consider two separate parameterisations of an intrinsic mass step with BayeSN and two slightly different samples; one based on the DES5YR cosmology sample after standard cuts and one using the full sample including redder SNe Ia. Neither of these choices substantially impacts our conclusions relating to dust. The largest $\Delta\mu_{R_{V}}$ we infer is for the full sample when allowing for an intrinsic SED difference on either side of the mass step. This analysis yields $\mu_{R,HM}=2.53\pm 0.15$ and $\mu_{R,LM}=3.25\pm 0.22$, with $\Delta\mu_{R_{V}}=-0.72\pm 0.26$. Previous analysis of the DES5YR sample using Dust2Dust yielded very different results, with best fit values of $\mu_{R,HM}=1.66$ and $\mu_{R,LM}=3.25$ with $\Delta\mu_{R_{V}}=1.59$.
Overall, we find that BayeSN is able to recover the dust properties from SALT-based simulations, but disagrees with the results from the DES5YR paper (Vincenzi et al., 2024). We propose the explanation that separately fitting the chromatic scatter component (e.g., as Vincenzi et al. (2024) did) from other observed features of SNe Ia will bias the inferred $R_{V}$ populations towards a larger $\Delta\mu_{R_{V}}$, by forcing the different $R_{V}$ populations to fully explain the observed mass step. A more recent analysis of the DES5YR sample, Popovic et al. (2024a), provides further evidence for this hypothesis: they find that the addition of a grey component to account for scatter and magnitude differences not only improves model fits, but also that a smaller $\Delta\mu_{R_{V}}$ (more consistent with the results of this paper) is needed when including non-chromatic scatter in the model fit.
@@ 7.1 The Impact of Dust on Photometric Classification
The $P_{\rm Ia}$ cut removes a portion of highly-reddened SNe Ia, truncating the tail of the exponential $E(B-V)$ distribution. In practice, the ‘effective’ $\tau_{E}$ value of the sample after cuts will be less than the true $\tau_{E}$ value. BayeSN is able to recover this effective value successfully on a sample with a $P_{\rm Ia}$ cut applied; however, this change in the simulated $\tau_{E}$ population indicates a previously-undiscovered need for improved photometric classification of highly-reddened SNe Ia, and a more streamlined integration of SN Ia classification into dust-inference pipelines.
@@ 7.2 Summary
In this work we have for the first time applied BayeSN to simulations based on SALT to test for consistency. We also apply BayeSN to a volume-limited subset of the DES5YR sample of SNe Ia to constrain the environmental-dependence of SN Ia host galaxy dust and intrinsic properties. Our key findings are summarised below:
  • We verify that the treatment of dust parameters between BayeSN and SALT is consistent, and demonstrate that BayeSN is able to accurately recover simulated dust properties from SALT-based simulations as used in the DES5YR analysis.
@@
  • We find that inferring dust parameters on a simulated DES5YR sample that contains non-Ia contamination is biased; this bias is not due to the choice of dust-inference method, but rather the inability of photometric classification to accurately identify the reddest, dimmest SNe Ia.
@@
  • We follow up G24 with ${>}3\sigma$ evidence to support an environmental-dependence of the secondary maximum of SNe Ia in the $i$-band when analysing a near-independent sample.
@@ Acknowledgements
MG and BP would like to thank Kaisey Mandel and Dan Scolnic for their valuable comments on the manuscript.

### v3

@@ 7 Discussion and Conclusions
Previous analyses of the host galaxy dust properties of SNe Ia using SALT and BayeSN have generally found conflicting results. SALT analyses have found large differences in the mean of the $R_{V}$ distribution, $\mu_{R_{V}}$, between SNe Ia on either side of the mass step, finding $\Delta\mu_{R_{V}}\sim 1.2-1.6$ (Popovic et al., 2023b; Vincenzi et al., 2024); this approach did not, until recently, allow for achromatic intrinsic differences (Popovic et al., 2024a). BayeSN, on the other hand, has supported the existence of an intrinsic contribution to the mass step and predicted $\Delta\mu_{R_{V}}$ values which are smaller (G24) or consistent with zero (Thorp et al., 2021; Thorp & Mandel, 2022).
In this work we have for the first time validated the performance of BayeSN by applying it for inference of dust properties on SALT-based simulations similar to those used in the DES5YR cosmology analysis. Despite the fundamental differences between the models, we find that BayeSN is able to accurately recover simulated dust properties from these SALT-based simulations for a wide variety of permutations of different simulations.
Having established this based on simulations, we then apply BayeSN to real data from the DES5YR sample. We consider two separate parameterisations of an intrinsic mass step with BayeSN and two slightly different samples; one based on the DES5YR cosmology sample after standard cuts and one using the full sample including redder SNe Ia. Neither of these choices substantially impacts our conclusions relating to dust. The largest $\Delta\mu_{R_{V}}$ we infer is for the full sample when allowing for an intrinsic SED difference on either side of the mass step. This analysis yields $\mu_{R,HM}=2.53\pm 0.15$ and $\mu_{R,LM}=3.25\pm 0.22$, with $\Delta\mu_{R_{V}}=-0.72\pm 0.26$. Previous analysis of the DES5YR sample using Dust2Dust yielded very different results, with best fit values of $\mu_{R,HM}=1.66$ and $\mu_{R,LM}=3.25$ with $\Delta\mu_{R_{V}}=-1.59$.
Overall, we find that BayeSN is able to recover the dust properties from SALT-based simulations, but disagrees with the results from the DES5YR paper (Vincenzi et al., 2024). We propose the explanation that separately fitting the chromatic scatter component (e.g., as in Vincenzi et al., 2024) from other observed features of SNe Ia will bias the inferred $R_{V}$ populations towards a larger $\Delta\mu_{R_{V}}$, by forcing the different $R_{V}$ populations to fully explain the observed mass step. A more recent analysis of the DES5YR sample, Popovic et al. (2024a), provides further evidence for this hypothesis: they find that the addition of a grey component to account for scatter and magnitude differences not only improves model fits, but also that a smaller $\Delta\mu_{R_{V}}$ (more consistent with the results of this paper) is needed when including non-chromatic scatter in the model fit.
@@
One potential cause of time-dependent differences between SNe Ia in different environments is a variable impact of dust during the phase range covered by the model (Förster et al., 2013). BayeSN assumes $A_{V}^{s}$ and $R_{V}^{s}$ are constant for each SN $s$ across the time period covered by the model. Previous studies, however, have suggested that dynamic, or even static, dust can scatter light from the supernova, leading to changes in the observed E(BV)E(B-V) and $R_{V}$ over the duration of the SN light curve (e.g. Bulla et al., 2018a, b), driven by time delays of the SN light entering the observer’s line of sight. This could potentially drive time-dependent differences between populations of SNe Ia in different environments. However, these SNe Ia with a time-varying impact of dust are typically found to be particularly red SNe such as SN 2014J (see Table 4 of Bulla et al. (2018b)). These would be entirely excluded from cosmology samples such as our Cosmo sample that nevertheless shows this time-dependent intrinsic $i$-band difference.
@@ 7.1 The Impact of Dust on Photometric Classification
The $P_{\rm Ia}$ cut removes a portion of highly-reddened SNe Ia, truncating the tail of the exponential E(BV)E(B-V) distribution. In practice, the ‘effective’ $\tau_{E}$ value of the sample after cuts will be less than the true $\tau_{E}$ value. BayeSN is able to recover this effective value successfully on a sample with a $P_{\rm Ia}$ cut applied; however, this change in the simulated $\tau_{E}$ population indicates a previously-undiscovered need for improved photometric classification of highly-reddened SNe Ia, and a more streamlined integration of SN Ia classification into dust-inference pipelines.
@@ 7.2 Summary
In this work we have for the first time applied BayeSN to simulations based on SALT to examine differences between the two approaches and validate the performance of BayeSN. We also apply BayeSN to a volume-limited subset of the DES5YR sample of SNe Ia to constrain the environmental-dependence of SN Ia host galaxy dust and intrinsic properties. Our key findings are summarised below:
  • We verify that BayeSN is able to accurately recover simulated dust properties from SALT-based simulations as used in the DES5YR analysis, despite the fundamental differences between the models.
  • We demonstrate that BayeSN is able to successfully disentangle differences in dust properties from intrinsic differences across the SN Ia population when applied to SALT-based simulations.
@@
  • We find that inferring dust parameters on a simulated DES5YR sample that contains non-Ia contamination is biased when using the photometric classifier applied for the real DES5YR sample. This bias arises because this classifier has reduced accuracy when applied to the reddest, dimmest SNe Ia rather than the choice of dust-inference method.
@@
  • We follow up G24 with $3\sigma$ evidence to support an environmental-dependence of the secondary maximum of SNe Ia in the $i$-band when analysing a near-independent sample.
@@ Acknowledgements
MG and BP would like to thank Kaisey Mandel, Dan Scolnic and Lisa Kelsey for their valuable comments on the manuscript.